White light containing wavelengths from 400 nm to 750 nm falls on a grating with 7500 lines/cm. How wide is the first-order spectrum on a screen 2.20 m away?
___m
Answer #1

For \theta1 for wavelength 400 nm, we have

Using an equation,

d sin \theta1 = m \lambda1                                                                          { eq.1 }

where, m = diffraction order = 1

\lambda1 = wavelength of white light = 400 x 10-9 m

d = diffraction grating = 1 / (7500 lines/cm)

inserting these values in above eq.

sin \theta1 = (1) (400 x 10-9 m) / [(0.01 m / 7500)]

sin \theta1 = (400 x 10-9 m) (7500) / (0.01 m)

\theta1 = sin-1 (0.3)

\theta1 = 17.460

For \theta2 for wavelength 750 nm, we have

sin \theta2 = (1) (750 x 10-9 m) / [(0.01 m / 7500)]

sin \theta2 = (750 x 10-9 m) (7500) / (0.01 m)

\theta2 = sin-1 (0.562)

\theta2 = 34.190

width of the first-order spectrum on a screen which will be given as :

w = D (tan \theta2 - tan \theta1)                                                                    { eq.2 }

where, D = separation distance = 2.2 m

inserting the values in eq.2,

w = (2.2 m) [tan 34.190 - tan 17.460]

w = (2.2 m) [(0.6793) - (0.3145)]

w = (2.2 m) (0.3648)

w = 0.8 m